OpenStudy (anonymous):

im meant to solve 2^x+3^y=z^2 and have got to 2^a=3^b-1, what does this tell me?

OpenStudy (anonymous):

oh and a=k+1 where x=2k

OpenStudy (kinggeorge):

Do you know anything about b?

OpenStudy (anonymous):

We now know x is even, let x=2k. So we have: 3^y=(z-2^k )(z+2^k ) Now (z-2^k ) and (z+2^k ) should both be powers of three, so therefore (z+2^k )-(z-2^k ) should equal a power of three, so 3^w=(z+2^k )-(z-2^k ) 3^w≠2^2k=2^x So 2^x should be divisible by three, but it isn’t. So let (z-2^k )=3^a and (z+2^k )=3^b then we have █((z+2^k )=3^b@(z-2^k )=3^a )/(2^(k+1)=3^b-3^a ) If a>0 then 3^b-3^a will be a power of three and 2^(k+1) is not divisible by three, so let a=0. We are left with 2^(k+1)=3^b-1

OpenStudy (kinggeorge):

Hold on. Why do we know x is even?

OpenStudy (anonymous):

So z^2 ≡1 (mod 3) unless z^2=3m for any integer m, and if x is odd, say x=(2k+1) for some integer k. However when z=3, we have 2^3+3^0=3^2, so one solution is when x=3,y=0 and z=3 When x is odd 2^(2k+1)≡2(mod 3) and when x is even 2^x≡1 (mod3) so z^2 is congrgent to 2^x when x is even.

OpenStudy (anonymous):

assuming that is correct

OpenStudy (anonymous):

What do you mean by "solve"?

OpenStudy (anonymous):

find solutions for x y and z where thsi equality holds

OpenStudy (anonymous):

In integers....

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

Number theory/Diophantine equations, right?

OpenStudy (anonymous):

yes i think so, apparently 2^a=3^b-1 is meant to tell me something or maybe represent something, but im not sure what

OpenStudy (anonymous):

Well it gives you a relation between a and b and therefore to k. This is a subfield of Diophantine Equations called Exponential Diophantine Equations and there is no general theory covering them, you just play it by ear, follow your nose, whatever.

OpenStudy (anonymous):

ok, never heard of that. any ideas on how i shoudl go about finishing it then?

OpenStudy (anonymous):

I didn't actually read all the arguments above but if there is aquery about odd/even, probably best to it by cases.

OpenStudy (kinggeorge):

It looks like you're trying to evaluate mod 3 only. It might also be helpful to valuate mod 2.

OpenStudy (anonymous):

what should i be evaluating at mod 2?

OpenStudy (kinggeorge):

Well, both sides of the equation. You can tell that \(z^2\) is even if and only if \(x=0\). So fix \(x=0\) and see if that gets you anywhere.

OpenStudy (anonymous):

i think i did that at the beginning, because z is odd unless x=0, but thas only one solution the solutions i have so far are X 0 4 3 Y 1 2 0 Z 2 5 3 however im not sure if there are more after these

OpenStudy (kinggeorge):

I tend to think those are the only ones, but I'm not sure how you would do that.

OpenStudy (anonymous):

yer that's the trouble im having haha